OpenStudy (anonymous):

one more like the previous Q 8! / 4! * 4! = 1 70 2!/4! 700

OpenStudy (anonymous):

70

OpenStudy (anonymous):

70

OpenStudy (anonymous):

40320/24^2=70

OpenStudy (anonymous):

ThankS!

OpenStudy (anonymous):

(8 x 7 x 6 x 5 x 4 x 3 x 2)/(4 x 3 x 2 x 4 x 3 x 2) = 70

OpenStudy (anonymous):

reasoning: \[\frac{8!}{4!4!} = \frac{5\times 6 \times 7 \times 8}{4!}\] since one 4! on the denominator cancels with the first four terms of 8! in the numerator. then just do \[\frac{5\times 7 \times 8}{4} = 280/4 = 70\]